OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form.

OpenStudy (anonymous):

Find the slope: \[\frac{y2-y1}{x2-x1}=\frac{-3-6}{0-3}=3\] You can then use point slope form to write the equation and switch to slope intercept form easily \[y-6=3(x-3), y=3x-9+6=3x-3\]